OpenStudy (anonymous):

Identify the slope and y-intercept of the following equation : 25=15x+5y and please show the solution .. Thanks . :)) , last one for the slope intercept form

OpenStudy (anonymous):

25=15x+5y -5y=15x-25 y=-3x+5 slope = -3 intercept (0,5)